OpenStudy (anonymous):

Write an equation for each line in point slope form and then convert it to standard form. Through (2,3) and (3,5)

OpenStudy (texaschic101):

first we find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (2,3)...x1 = 2 and y1 = 3 (3,5)...x2 = 3 and y2 = 5 now we sub slope(m) = (5 - 3) / (3 - 2) slope(m) = 2/1 or just 2 point slope form : y - y1 = m(x - x1) slope(m) = 2 you can use either of your points..(2,3)...x1 = 2 and y1 = 3 now we sub y - 3 = 2(x - 2) <=== point slope form y - 3 = 2x - 4 y = 2x - 4 + 3 y = 2x - 1 <===slope intercept form y = 2x - 1 --- subtract 2x from both sides -2x + y = -1 -- multiply by -1 2x - y = 1 <=== standard form or you could have used point slope form with the other point (3,5)...x1 = 3 and y1 = 5 slope(m) = 2 y - 5 = 2(x - 3) <== point slope form y - 5 = 2x - 6 y = 2x - 6 + 5 y = 2x - 1 <== slope intercept form and you can turn it into standard form the same way I did before and you should get the same answer in standard form your point slope forms for these 2 sets of points may be different, but the standard form will be the same

OpenStudy (anonymous):

Got it. Do you mind helping me with one more? @texaschic101